Using power series method to solve the initial value problem

(x − )1 y′′ + yx ′ + y = 0 , y )0( = 2 , y′ )0( = −1 .


1
Expert's answer
2021-06-03T15:32:46-0400

given, the differential equation

(x − 1) y′′ +x y′ + y = 0,

y (0)= 2 , y′(0)= −1.

Let the solution be "y=\sum c_nx^n \newline
y'=\sum nc_nx^{n-1} \newline
y''=\sum n(n-1)c_nx^{n-2} \newline
\text{substitute the above value in the given equation, we get}\newline
(x-1)\sum_{n=2}^\infty n(n-1)c_nx^{n-2}+x\sum_{n=1}^\infty nc_nx^{n-1}+\sum_{n=0}^\infty c_nx^n=0 \newline
\sum_{n=2}^\infty n(n-1)c_nx^{n-1}-\sum_{n=2}^\infty n(n-1)c_nx^{n-2}
-\sum_{n=1}^\infty nc_nx^{n}+\sum_{n=0}^\infty c_nx^n=0 \newline
\sum_{n=2}^\infty n(n+1)c_{n+1}x^{n}-\sum_{n=2}^\infty (n+1)(n+2)c_{n+2}x^{n}
-\sum_{n=1}^\infty nc_nx^{n}+\sum_{n=0}^\infty c_nx^n=0 \newline
2c_2-c_1+\sum_{n=0}^\infty n(n+1)c_{n+1}x^{n}-\sum_{n=0}^\infty (n+1)(n+2)c_{n+2}x^{n}
-\sum_{n=0}^\infty nc_nx^{n}+\sum_{n=0}^\infty c_nx^n=0 \newline
\text{We get,}\newline
2c_2-c_1=0
\text{and
the recurrence relation is,}
c_{n+2}=\frac{n(n+1)c_{n+1}+(1-n)c_{n}}{(n+1)(n+2)}.\newline
\text{Thus, the solution is given by }\newline
y=c_0y1+c_1y_2.\\
Let c_0=1, c_1=0, \implies y=y_1.\\
Then, c_2=1/2.\\
Therefore, y_1=1+\frac{1}{2}x^2+... .\\
Similarly,\newline
 c_0=0, c_1=1, \implies y=y_2.\newline
Then, c_2=0, c_3=0.\newline
Therefore,
y_2=x"